OpenStudy (firejay5):

Describe the change to the parent function. *I don't really get these at all much* 1. y = -3x^2 2. y = 2/5x^2 3. y = x^2 + 3 4. y = x^2 - 2 5. y = (x + 2) ^2 6. y = (x - 4)^2 7. y = 2x^2 + 1 8. y = -x^2 - 5 9. y = 4(x + 2)^2 - 1 10. y = -1/3(x - 1)^2 + 2

OpenStudy (firejay5):

My Answers: Choices (reflection over x or y - axis, hz or vt stretch, hz or vt compression, up/down, left, right) 1. reflection over x-axis and vt stretch 2. vt stretch 3. up 3 units 4.down 2 units 5. left 2; 6. right 4; 7. up 1 unit; 8. down 5 units; 9. left 2 units; down 1 unit; vt. compression 10. up 2 units;

OpenStudy (firejay5):

@jim_thompson5910 @Loser66 @Directrix @dtan5457 @Luigi0210 @jagr2713 Help me please, I have the answers and I need to make sure if there is anything I need to add to any questions

jimthompson5910 (jim_thompson5910):

1. vertical stretch by what factor? 2. vertical stretch by what factor? 3. correct 4. correct 5. correct 6. correct 7. you're missing the vertical stretch 8. you have a reflection going on here 9. vertical stretch by what factor? 10. you forgot about the vertical stretch and horizontal translation

OpenStudy (firejay5):

@jim_thompson5910 what do you mean by vt stretch by what factor; #8 is reflection over x -axis; what do you mean by hz translation on #10

jimthompson5910 (jim_thompson5910):

example: parent function is y = x^2 the function turns into y = 9x^2 the vertical stretch is by a factor of 9 (so each point's y coord is multiplied by 9, ie it's 9 times taller in a way)

jimthompson5910 (jim_thompson5910):

y = -1/3(x - 1)^2 + 2 has a horizontal translation of 1 to the right (we shift 1 unit to the right) due to that x-1

OpenStudy (firejay5):

so #10 is a vt stretch of -1/3

jimthompson5910 (jim_thompson5910):

also, 10 has a reflection (because of the negative 1/3)

OpenStudy (firejay5):

reflection over x-axis

jimthompson5910 (jim_thompson5910):

I'd say "vertical stretch of 1/3" not -1/3 you usually say the magnitude of how much you're stretching

OpenStudy (firejay5):

so am I good on my answers @jim_thompson5910: My Answers: Choices (reflection over x or y - axis, hz or vt stretch, hz or vt compression, up/down, left, right) 1. reflection over x-axis and vt stretch of 3 2. vt stretch of 2/5 3. up 3 units 4.down 2 units 5. left 2 6. right 4 7. up 1 unit; vt stretch of 2 8. down 5 units; reflection over x-axis 9. left 2 units; down 1 unit; vt. compression 10. up 2 units; vt stretch of 1/3; shifted 1 unit to the right; reflection over x-axis

jimthompson5910 (jim_thompson5910):

everything looks good, but 9 is still missing the amount you're stretching vertically

jimthompson5910 (jim_thompson5910):

and because the '4' is larger than 1, you aren't compressing you're stretching

jimthompson5910 (jim_thompson5910):

# 2 and #10 are vertical compressions

OpenStudy (firejay5):

I thought we said 2 and 10 were vertical stretches

jimthompson5910 (jim_thompson5910):

notice the 2/5 is less than 1 that means we have a compression it's a dilation and you can refer to it like that if you want

jimthompson5910 (jim_thompson5910):

read this article http://www.regentsprep.org/regents/math/algtrig/atp9/funclesson1.htm

OpenStudy (firejay5):

so #9 is vt compression of 4

jimthompson5910 (jim_thompson5910):

focus on the "Stretch or Compress Functions" section

jimthompson5910 (jim_thompson5910):

no the 4 outside is larger than 1

jimthompson5910 (jim_thompson5910):

y = a(x-h)^2 + k if a > 1 or a < -1, then we have vertical stretching if -1 < a < 1, then we have vertical compression

OpenStudy (firejay5):

so #9 is vt stretch of 4 then

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (firejay5):

so is all of my answers in good top shape now lol

jimthompson5910 (jim_thompson5910):

yes they look good

OpenStudy (firejay5):

@jim_thompson5910 I was wondering how #1 isn't a vertical compression since its less than 1?

jimthompson5910 (jim_thompson5910):

y = a(x-h)^2 + k if a > 1 or a < -1, then we have vertical stretching if -1 < a < 1, then we have vertical compression

jimthompson5910 (jim_thompson5910):

a = -3 fits the first description: "if a > 1 or a < -1, then we have vertical stretching"

jimthompson5910 (jim_thompson5910):

read this page http://www.regentsprep.org/regents/math/algtrig/atp9/funclesson1.htm read everything on it, but pay special attention to the "Stretch or Compress Functions" section